OpenStudy (anonymous):

The sum of the 32 terms in an arithmetic series is 5488 and the first term is 1. Find the last term in the series.

OpenStudy (anonymous):

5488 = 32 * (1 + x)/2 thus 1 + x = 5488/16 = 343 the answer is x = 342

OpenStudy (anonymous):

is der any doubt?

OpenStudy (anonymous):